OpenStudy (anonymous):

Use scientific notation to multiply the following two numbers. Express the answer using scientific notation; retain all nonzero decimal places. (8.1 · 10^7) · (4.2 · 10^9)

Parth (parthkohli):

81000000 * 4200000000 = 340200000000000000 =3.402 * 10^16

OpenStudy (anonymous):

\[8,1*4,2*10^{7+9}=34,02*10^{16}\]

OpenStudy (anonymous):

thanks